# How much has DESI dark energy evolved since DR1?

### v1

@@ II Analysis
In summary, our analysis exposes a $q_{0}$ sign problem and evident fluctuations in DESI DR2 BAO. As is clear from Fig. 4, LRG3+ELG1 shows excellent agreement between DR1 and DR2, while ELG2 and Lyman-$\alpha$ QSO also show good agreement. On the other hand, LRG1 has become more consistent with Planck. The most interesting difference between DR1 and DR2 is the shift in LRG2, which means it goes from contributing to the lower $\Omega_{m}$ preferred by the full sample relative to Planck, to the tracer that is most sharply driving the dynamical DE signal. In Fig. 5 we remind the reader that despite the potential for statistical fluctuations in DESI BAO data, when DESI DR1 BAO is combined with FS modeling, there are no obvious outliers and all constraints intersect the $\Omega_{m}$ value for the full BAO+FS dataset Adame et al. (2024b, c). As a result, there is no signal of dynamical DE in DESI DR1 data alone. That being said, DESI DR1 data prefers a value for the $\Lambda$CDM parameter $\Omega_{m}$ that is $1.6\sigma$ lower than the Planck value. Admittedly, this constant shift in $\Omega_{m}$ challenges concordance, but claims that $\Omega_{m}$ is not a constant in $\Lambda$CDM cosmology, and is in fact redshift dependent, go back to 2022 Ó Colgáin et al. (2022, 2024). The main point here is that a non-constant $\Omega_{m}$ $\Lambda$CDM parameter, while pointing to model breakdown, does not immediately imply a dynamical DE sector.
@@
Table 3: Constraints on the CPL model from the full sample with LRG1 and LRG2 data removed.
Figure 4: Differences in the $\Lambda$CDM $\Omega_{m}$ constraints from individual tracers between DR1 in faded blue and DR2 in blue. The red and green bands denote Planck and full DESI DR2 BAO sample constraints on $\Omega_{m}$. In magenta we separate LRG3 and ELG1 constraints.
Figure 5: DESI DR1 BAO constraints on the $\Lambda$CDM parameter $\Omega_{m}$ in faded blue relative to the DESI DR1 BAO+FS modeling constraints in blue. The green strip denotes the constraint on the full sample from BAO+FS and it can be confirmed that all constraints show excellent agreement.
@@ III Discussion
Any dynamical DE signal with $w_{0}>-1$ cannot be the final word on a replacement for the $\Lambda$CDM model because the $w_{0}-H_{0}$ anti-correlation is problematic in the face of $H_{0}>70$ km/s/Mpc determinations Vagnozzi et al. (2018); Vagnozzi (2020); Alestas et al. (2020); Banerjee et al. (2021); Lee et al. (2022). A key point is that larger than expected local $H_{0}$ values are observed across multiple observables. As a result, one infers that there must be something wrong with the DESI dynamical DE claim Adame et al. (2024a, b); Abdul Karim et al. (2025), at least in its current form. It is true that the CPL model fits DESI+CMB+SNe and DESI+CMB datasets better than $\Lambda$CDM in a statistically significant manner, but if it contradicts Hubble tension, or worse, fails to confirm late-time accelerated expansion today, one should bin the idea. That being said, it should be borne in mind that Hubble tension is an expected harbinger new physics beyond $\Lambda$CDM.
In this letter we looked at the improvements in DESI BAO data between DR1 Adame et al. (2024a) and DR2 Abdul Karim et al. (2025). When the data is confronted to the CPL model, we noted that the narrow $w_{a}\in[-3,2]$ priors employed by the DESI collaboration cut off the $w_{a}$ posterior, giving rise to skewed posteriors that become more symmetric once the priors are relaxed. In agreement with earlier work Wang (2024), with relaxed priors we confirmed that DESI DR1 BAO is inconsistent with late-time accelerated expansion today ($q_{0}<0$). In the upgrade to DESI DR2 BAO, we note that $q_{0}<0$ cannot be confirmed, even when one combines DR2 BAO with CMB. Thus, we have a $3.1\sigma$ deviation from $\Lambda$CDM Abdul Karim et al. (2025), yet cannot confirm $q_{0}<0$.
@@
We observe that LRG1, which was the most prominent outlier in DR1 BAO Ó Colgáin et al. (2024), is now consistent with a lower $\Omega_{m}$ value in the $\Lambda$CDM model. It is still the most prominent outlier at $1.8\sigma$ removed from the rest of the dataset. Moreover, we find that LRG2 returns an $\Omega_{m}$ value larger than Planck in DR2 in contrast to the smaller value in DR1. We now confirm that ELG data is solely responsible for the lower $\Omega_{m}$ relative to Planck in the full DR2 BAO sample. Finally, we show that LRG2 and not LRG1 is now most responsible for the $w_{0}>-1$ dynamical DE signal.
Evidently, fluctuations still persist in DESI BAO data and we have yet to see convergence in constraints for both LRG1 and LRG2, although high redshift tracers show good to excellent agreement between DR1 and DR2. It will be interesting to see if any fluctuations remain when DR2 BAO is combined with FS modeling, as there is no dynamical DE signal when DR1 BAO is combined with DR1 FS modeling Adame et al. (2024c).
@@ Acknowledgements
We would like to thank Eleonora Di Valentino and Sunny Vagnozzi for correspondence on credible intervals. This article/publication is based upon work from COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”, supported by COST (European Cooperation in Science and Technology).

### v2

@@ II Analysis
In summary, our analysis exposes a $q_{0}$ sign problem and evident fluctuations in DESI DR2 BAO. As is clear from Fig. 4, LRG3+ELG1 shows excellent agreement between DR1 and DR2, while ELG2 and Lyman-$\alpha$ QSO also show good agreement. On the other hand, LRG1 has become more consistent with Planck. LRG1 and ELG1 are outliers at $1.8\sigma$ with respect to the value of $\Omega_{m}$ preferred by the full DR2 BAO sample. However, the most interesting difference between DR1 and DR2 is the shift in LRG2, which means it goes from contributing to the lower $\Omega_{m}$ preferred by the full sample relative to Planck, to the tracer that is most sharply driving the dynamical DE signal (see Fig. 4). In Fig. 5 we remind the reader that despite the potential for statistical fluctuations in DESI BAO data, when DESI DR1 BAO is combined with FS modeling, there are no obvious outliers and all constraints intersect the $\Omega_{m}$ value for the full BAO+FS dataset Adame et al. (2024b, c), i. e. the equivalent of LRG1 and ELG1 BAO constraints do not exist. As a result, there is no signal of dynamical DE in DESI DR1 data alone. That being said, DESI DR1 BAO+FS data prefers a value for the $\Lambda$CDM parameter $\Omega_{m}$ that is $1.6\sigma$ lower than the Planck value. Admittedly, this constant shift in $\Omega_{m}$ challenges concordance, but claims that $\Omega_{m}$ is not a constant in $\Lambda$CDM cosmology, and is in fact redshift dependent, go back to 2022 Ó Colgáin et al. (2022, 2024). The main point here is that a non-constant $\Omega_{m}$ $\Lambda$CDM parameter, while pointing to model breakdown, does not immediately imply a dynamical DE sector.
@@
Table 3: Constraints on the CPL model from the full sample with LRG1, ELG1 and LRG2 data removed.
Figure 4: Differences in the $\Lambda$CDM $\Omega_{m}$ constraints from individual tracers between DR1 in faded blue and DR2 in blue. The red and green bands denote Planck and full DESI DR2 BAO sample constraints on $\Omega_{m}$. In magenta we separate LRG3 and ELG1 constraints. We have displaced redshifts for visually purposes.
Figure 5: DESI DR1 BAO constraints on the $\Lambda$CDM parameter $\Omega_{m}$ in faded blue relative to the DESI DR1 BAO+FS modeling constraints in blue. The green strip denotes the constraint on the full sample from BAO+FS and it can be confirmed that all constraints show excellent agreement. Ideally, BAO and FS modeling should agree on $\Omega_{m}$.
@@ III Discussion
Any dynamical DE signal with $w_{0}>-1$ cannot be the final word on a replacement for the $\Lambda$CDM model because the $w_{0}-H_{0}$ anti-correlation is problematic in the face of $H_{0}>70$ km/s/Mpc determinations Vagnozzi et al. (2018); Vagnozzi (2020); Alestas et al. (2020); Banerjee et al. (2021); Lee et al. (2022). A key point is that larger than expected local $H_{0}$ values are observed across multiple observables. As a result, one infers that there must be something wrong with the DESI dynamical DE claim Adame et al. (2024a, b); Abdul Karim et al. (2025), at least in its current form. It is true that the CPL model fits DESI+CMB+SNe and DESI+CMB datasets better than $\Lambda$CDM in a statistically significant manner, but if it contradicts Hubble tension, or worse, fails to confirm late-time accelerated expansion today, one should rethink. [^5] Physics traditionally makes progress by identifying and resolving contradictions. That being said, it should be borne in mind that Hubble tension is an expected harbinger of new physics beyond $\Lambda$CDM.
In this letter we looked at the improvements in DESI BAO data between DR1 Adame et al. (2024a) and DR2 Abdul Karim et al. (2025). When the data is confronted to the CPL model, we noted that the narrow $w_{a}\in[-3,2]$ priors employed by the DESI collaboration cut off the $w_{a}$ posterior, giving rise to skewed posteriors that become more symmetric once the priors are relaxed. In agreement with earlier work Wang (2024), with relaxed priors we confirmed that DESI DR1 BAO is inconsistent with late-time accelerated expansion today ($q_{0}<0$). In the upgrade to DESI DR2 BAO, we note that $q_{0}<0$ cannot be confirmed, even when one combines DR2 BAO with CMB. Thus, we have a $3.1\sigma$ deviation from $\Lambda$CDM Abdul Karim et al. (2025), [^6] Note, DESI DR2 BAO and CMB have an approximate $1.6\sigma$ discrepancy in the $\Lambda$CDM parameters $\Omega_{m}$. This can only translate into a $3.1\sigma$ deviation from $\Lambda$CDM in the CPL model Abdul Karim et al. (2025) if BAO data has an inherent dynamical DE signal. Replacing BAO with FS modeling, where there is no inherent dynamical DE signal, the statistical significance of the deviation must drop. yet cannot confirm $q_{0}<0$. This may have tantalising implications for theoretical speculation Brandenberger (2019, 2025); Li et al. (2025), but more mundanely may simply highlight a problem fitting the CPL model to datasets with a high effective redshift.
@@
We observe that LRG1, which was the most prominent outlier in DR1 BAO Ó Colgáin et al. (2024), is now consistent with a lower $\Omega_{m}$ value in the $\Lambda$CDM model. It is still jointly the most prominent outlier at $1.8\sigma$ removed from the rest of the dataset. In addition, we see that ELG1 prefers an $\Omega_{m}$ value $1.8\sigma$ lower than the full sample. Moreover, we find that LRG2 returns an $\Omega_{m}$ value larger than Planck in DR2 in contrast to the smaller value in DR1. We now confirm that ELG data is solely responsible for the lower $\Omega_{m}$ relative to Planck in the full DR2 BAO sample. Finally, we show that LRG2 and neither LRG1 nor ELG1 is now most responsible for the $w_{0}>-1$ dynamical DE signal.
Evidently, fluctuations still persist in DESI BAO data and we have yet to see convergence in constraints for both LRG1 and LRG2, although high redshift tracers show good to excellent agreement between DR1 and DR2. In addition, ELG1 results in an unusually low $\Omega_{m}$ value. It will be interesting to see if any fluctuations remain when DR2 BAO is combined with DR2 FS modeling, as there is no dynamical DE signal when DR1 BAO is combined with DR1 FS modeling Adame et al. (2024c).
@@ Acknowledgements
We would like to thank Eleonora Di Valentino and Sunny Vagnozzi for correspondence on credible intervals. This article/publication is based upon work from COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”, supported by COST (European Cooperation in Science and Technology). L. Yin was supported by Natural Science Foundation of Shanghai 24ZR1424600.
Appendix A Simple argument on $w(z)$ and $H_{0}$
In Vagnozzi et al. (2018); Vagnozzi (2020); Alestas et al. (2020); Banerjee et al. (2021); Lee et al. (2022) it has been shown pretty generally that increasing $w(z)=-1$ ($\Lambda$CDM) to $w(z)>-1$ (quintessence) lowers $H_{0}$. Here is an analytic argument. In general, the DE density $X(z)$ is related to the equation of state $w(z)$ through:
$X(z)=\exp\left(3\int_{0}^{z}\frac{1+w(z^{\prime})}{1+z^{\prime}}\textrm{d}z^{\prime}\right).$ (3)
Observe that $w(z)>-1\Rightarrow X(z)>1$ and for $\Lambda$CDM $X(z)=1$. Consider an $X(z)$CDM dynamical DE model with Hubble parameter
$H(z)^{2}=H_{0}^{2}(1-\Omega_{m})X(z)+H_{0}^{2}\Omega_{m}(1+z)^{3},$ (4)
where we restrict our attention to the late Universe. By definition the DE term is not supposed to be relevant at high redshift, as is the case in $\Lambda$CDM, hence the combination $H_{0}^{2}\Omega_{m}$ is expected to be essentially constrained the same in $\Lambda$CDM and in any replacement $X(z)$CDM model. It is easy to see now that if $H_{0}^{2}\Omega_{m}$ is a constant, but one increases $X(z)$ through a theoretical choice, then $H_{0}^{2}(1-\Omega_{m})$ with constant $H_{0}^{2}\Omega_{m}$, and hence $H_{0}$, must decrease to find a fit no worse than $\Lambda$CDM to the data. Note, the $X(z)$CDM model with $X(z)>1$ can fit the data better than $\Lambda$CDM, but the price one pays is a lower $H_{0}$. This guarantees that one cannot resolve $H_{0}$ tension, and errors aside, most likely makes the problem worse.
Our argument here covers analysis in Banerjee et al. (2021); Lee et al. (2022). In Banerjee et al. (2021) observed data was studied, and while quintessence models that raised $H_{0}$ were found, they all fitted the data worse than $\Lambda$CDM. Despite CMB data not being utilised, one infers that the combination $H_{0}^{2}\Omega_{m}$ must have remained constant between $\Lambda$CDM and quintessence fits to a combination of BAO, cosmic chronometer and SNe data. In contrast, Lee et al. (2022) worked with mock DESI forecast data up to higher redshifts $z\leq 3.55$. We emphasize that in both these exercises no exceptions were found; an increase in $w(z)$ always led to a decrease in $H_{0}$. As can be seen from the analytic argument, this is guaranteed provided both $\Lambda$CDM and the replacement DE model with $w(z)>-1$ lead to the same constraints on $H_{0}^{2}\Omega_{m}$. The non-constancy of $H_{0}^{2}\Omega_{m}$ is the only loophole.